# CARTOS: A Charging-Aware Real-Time Operating System for Intermittent Batteryless Devices

### v2

@@ Abstract
Index Terms:
Intermittent computing, scheduling, real-time and embedded systems, operating systems.
@@ I Introduction
BATTERYLESS systems, often called intermittently-powered devices (IPDs), have the potential to revolutionize a variety of applications in Internet-of-Things (IoT), such as smart healthcare, agriculture, and building monitoring, where the use of batteries is undesirable or incurs significant maintenance costs. These systems harvest energy from ambient sources such as sunlight, heat, and radio signals, charge a small energy buffer, and execute intermittently for a short period of time whenever the energy is available. Since no batteries are needed, they have a much longer lifespan than battery-powered systems and can be deployed in harsh environments where batteries cannot operate properly, e.g., due to cold and hot temperatures [1, 2, 3]. This can also bring other benefits to system design such as smaller form factors and reduced manufacturing costs.
The aforementioned benefits, however, come with several technical challenges. First, given the variability in energy availability from ambient sources, IPDs must ensure the forward progress of program execution across power failures. One of the most effective strategies is checkpointing [4, 5, 6, 7], a method that involves inserting additional code snippets to store the state of a running program in non-volatile memory (NVM). When power is restored, the system can recover from the latest stored state, ensuring continuous progress. However, such static checkpointing introduces a high overhead and can significantly impact the performance of intermittent systems compared to conventional battery-powered embedded systems. Just-In-Time (JIT) checkpointing [8, 9, 10] mitigates this issue by making checkpoints dynamically only when the power goes low. However, it is difficult to be used for tasks with peripheral operations, e.g., sensor access and flash writing, because they cannot be paused and resumed at an arbitrary point, i.e., execution needs to be done atomically.
Secondly, IPDs should provide accurate time behavior even in the presence of intermittent power disruptions. This is particularly important for periodic sensing applications, which are one of the most in-demand areas of IoT. While many battery-powered IoT devices have been developed using conventional real-time operating systems (RTOS) for microcontrollers, such as FreeRTOS, $\mu$C/OS, RTEMS, to implement timely operations, they cannot offer forward progress guarantees in IPDs. For instance, a power failure during the execution of a sensor-reading or data-logging task can result in incomplete task execution or the loss or corruption of data. Therefore, it is crucial for the RTOS to consider the power consumption of tasks and the energy availability of the system. Furthermore, providing analytical support to understand how these devices will operate under diverse real-world circumstances is of great importance.
@@
CARTOS is implemented by introducing new APIs and components to FreeRTOS, which is one of the most widely used embedded RTOS in real-world applications such as Amazon AWS IoT. However, as the design of CARTOS is not specific to FreeRTOS, it is portable to other RTOS as well. From our experiments on real hardware and simulation, CARTOS demonstrated several advantages over the state-of-the-art methods.
Section II provides an overview of previous work and establishes essential background information. In Section III, we introduce the proposed framework. Section IV presents the system model, the proposed mixed-preemption scheduling approach, and offers a detailed scheduling analysis. Section V presents the results of our evaluation. Lastly, Section VI summarizes and concludes this paper.
@@ II-A IPD Hardware and Operation Phases
A typical hardware design of IPDs is illustrated in Fig. 1. The harvesting unit collects energy from the surrounding environment. It could be a solar cell, RFID receiver, or piezoelectric transducer. The power management unit (PMU) is responsible for converting and regulating the energy harvested by the energy harvester into a stable voltage that can be used to power the device’s components. It also manages the device’s power consumption to ensure that it does not exceed the available energy. Energy storage is usually a supercapacitor and is responsible for storing the collected electrical energy so that it can be used later when it is needed. The micro-controller unit (MCU) executes the device’s software instructions and controls the device’s various components. Memory stores the program code and data that the MCU uses to operate the device. Most IPDs have at least two types of memory: (i) volatile main memory equipped as part of MCU, e.g., SRAM, and (ii) nonvolatile memory to store execution progress during power loss, e.g., FRAM and MRAM. IPDs also often include sensors and peripherals, such as temperature, humidity, light sensors and wireless Bluetooth communication modules, to monitor the device’s environment, perform certain actions, and report any findings to the user.
@@
    Operation Phase: Once the capacitor attains sufficient voltage, the PMU provides a consistent supply to the device. This enables the MCU to execute tasks. The device continues in this phase until it reaches the predefined power-off threshold (returning to the power-off phase) or low-voltage threshold (moving to the standby phase discussed next) depending on the design. Many prior IPDs [11, 12, 4] use only the power-off threshold for simplicity and operate as in Fig. 2. The MCU starts computation and sensing tasks when the capacitor voltage reaches a specific power-on threshold and continues until the voltage falls below the power-off threshold. Charging can still occur while the system runs tasks in this phase; however, as the discharging rate typically exceeds the charging rate, the capacitor voltage decreases over time.
@@
    Standby Phase: In this phase, energy-intensive components like the MCU, sensors, and radio units are either powered down or transitioned into sleep mode. Only the timekeeping ability, e.g., MCU’s clock unit or external programmable RTC, remains active as it plays a crucial role in maintaining the precise time behavior of tasks [13, 14, 15, 16]. The device remains in this state until the capacitor accumulates enough charge to return to the operation phase. This approach can be particularly useful when the energy-harvesting rate is sporadic, as it allows the device to conserve energy and maintain the notion of time during low power availability.
@@ II-B Forward Progress Mechanisms
There has been extensive research on checkpointing to ensure that programs running on IPDs make forward progress and maintain data consistency across power failures. Static checkpointing [4, 5, 6, 17, 18, 19] inserts checkpointing code at the fixed locations of the program, which are determined by either the compiler support or manual programmer efforts. When a checkpoint is reached, the system saves the program’s state and relevant data to non-volatile storage, allowing the program to restart from that point in case of a power failure. Instead of finding checkpoint locations after writing the program code, new programming models have been studied to create checkpoints as part of the program’s execution flow [6, 20, 21, 22, 7, 11, 12]. [^1] In the context of IPDs, these approaches are often referred to as task-based intermittent systems. However, it is important to note that “tasks” in these systems differ from our work: they represent short non-preemptible atomic blocks, not-preemptible threads as in conventional RTOS and our work. In such models, the programmer needs to decompose a program logic into a collection of idempotent, atomic blocks with explicit input and output data. Then, the input and output data of each atomic block are recorded in NVM, which serves as checkpoints, e.g., data channels between blocks [11, 7]. However, this adds a significant burden to programmers to restructure their programs, which is particularly difficult for long-running computations, e.g., algorithm logic. In addition, programmers have to make sure that each atomic block is small enough to run within one charging cycle.
Just-In-Time (JIT) checkpointing [10, 8, 23, 9], on the other hand, makes checkpoints at runtime by capturing the current state of the system, including MCU register values and volatile data. To do so, the system needs to monitor the remaining energy and initiate JIT checkpointing when the low-voltage threshold is reached. This can bring performance benefits over static checkpointing by reducing the frequency of checkpointing and simplifying the programming of long-running computations [23, 8]. However, since peripheral operations cannot resume after reboot by recovering just the MCU’s state, JIT checkpointing cannot be used for them [10].
@@
Mementos[6]
Chinchilla[4]
Capybara[21]
Quickrecall[8]
Samoyed[10]
InK[20]
CatNap[9]
Celebi[14]
Rtag[13]
Karimi et al.[24]
@@ II-C IPD Task Scheduling
Task scheduling for batteryless devices can be categorized into two main types: real-time and reactive scheduling. Real-time scheduling focuses on executing periodic tasks with predictable charging behavior. These approaches aim to meet the deadlines of a specific set of tasks by assuming a certain amount of energy supply[13, 24, 14]. By doing so, they can guarantee tasks to be scheduled in a predictable manner and achieve consistent performance as long as the assumed energy model holds. On the other hand, reactive scheduling focuses on minimizing the overall response times of event-driven tasks [9, 20]. It starts task execution when the device’s energy reaches the predefined power-on threshold and stops when it drops below the power-off threshold. These approaches are usually more flexible and may perform better with changing environmental conditions than the real-time approaches, but cannot ensure predictability and consistent performance due to their best-effort nature.
@@ II-D Related Work
Among static checkpointing approaches, Mementos [6] proposes a method to automatically insert checkpointing code to the program at compile time. To reduce the cost of checkpointing, it also proposes to estimate the remaining energy and perform checkpointing only when the energy goes below a certain threshold. Chinchilla [4] also takes a similar approach. Capybara [21] relies on the programming model support for forward progress, and introduces a hardware-software co-design solution that includes a set of capacitors to handle atomic blocks with different energy requirements. To address the limitations of programming model-based approaches that require the user to divide program logic into a series of atomic blocks, Immortal Threads [25] takes a compiler-assisted approach that transforms conventional threaded programs by inserting micro checkpoints. This allows the user to write programs in a typical multithreaded manner and enjoy preemption at the boundaries of checkpoints.
@@
Quickrecall [8] is a JIT checkpointing system. It detect low-voltage threshold and checkpoints preemptible tasks so that they can be resumed in next power cycle. However, it does not support peripheral tasks and restarts them in next power cycle if the power goes down during their execution. Samoyed [10] addresses this issue by introducing the undo-logging mechanism, which basically adds static checkpoints for peripheral operations. However, this approach is applicable only to the peripherals that “do not have internal non-volatile state and are arbitrarily restartable” [10] because of the idempotency requirement.
@@
Several approaches have been proposed to support peripheral devices in IPDs. Sytare [26] allows users to write additional code to explicitly store peripheral states in NVM. RESTOP [27] retains peripheral states by recording data transmission between MCU and peripherals, which incurs high overhead. EaseIO [28] takes a programming model-based approach that allows users to annotate I/O operations and transforms the program code based on this information. These methods offer various trade-offs between programmer effort, system overhead, and generality of peripheral support.
@@
For real-time scheduling on IPDs, Celebi [14] proposes a scheduler that optimizes task scheduling based on known charging patterns. However, this method is primarily designed for independent computational tasks and may not handle peripheral tasks correctly because the scheduler does not ensure the atomicity and idempotency of task execution, e.g., the scheduler can stop a task at an arbitrary point. Karimi at al. [13] introduce a real-time scheduler designed to atomically execute independent periodic tasks with various energy requirements and provide a formal analysis of task schedulability for a fixed charging rate of the system. The same authors [24] present an energy-adaptive scheduler to minimize the age of information. While these approaches contribute to enabling real-time operations on IPDs, none of them are sufficient to handle real-world applications that consist of a chain of peripheral access and long-running computations.
@@
There are also other previous studies that focus on intermittent neural networks [29, 15, 30], multiple energy buffers [31], energy-aware memory mapping [32], communication [33], design space exploration [34], and zero-power timekeeping [35]. They contribute to expanding the application scope of IPDs and our work does not conflict with them.
@@ III-A Hardware Setup
CARTOS targets an IPD that is equipped with an MCU capable of running a conventional RTOS such as FreeRTOS. The IPD is expected to have both volatile memory (SRAM) and non-volatile memory (NVM). Although we used Ambiq Apollo4 Evaluation board [36] which is equipped with an ARM Cortex-M4 MCU in our implementation, there is no other restriction imposed on MCUs by CARTOS.
@@ III-B Task Chains and Mixed Preemption Model
CARTOS supports periodic applications that consist of a processing chain of tasks, which is a common design pattern found in sensing and intelligent applications [37, 38, 11, 39, 40]. Fig 4 illustrates two processing chains, one comprising four tasks and the other with three tasks. Each task $\tau_{i}^{j}$ on a chain $i$ can be either a peripheral or computation task. [^2] Detailed chain and task parameters needed for analysis will be provided in Sec. IV. The chain can start execution when its first task is released (e.g., the periodic sensing task $\tau_{1}^{1}$ on chain 1) and the subsequent tasks are eligible to run only when their preceding task finishes (e.g., the action task $\tau_{1}^{4}$ becomes “ready” after the completion of the processing task $\tau_{1}^{3}$). We assume that the first task of a chain is a periodic task, triggered by either a timer or an external event. Hence, all other tasks of the same chain follow the same periodic pattern with precedence constraints among them. Chains have deadlines to meet. We consider constrained deadlines (deadline $\leq$ period) due to their wide acceptance. A formal model of tasks and chains will follow in Sec. IV
@@ III-B1 Mixed-Preemption Model
Recall that processing chains can consist of a mixture of peripheral and computation tasks. Peripheral tasks need to be executed atomically, especially when the corresponding hardware maintains internal state [10]. To ensure the atomic execution of peripheral tasks, CARTOS schedules peripheral tasks non-preemptively, i.e., once selected for scheduling, it will continue with no preemption from other higher-priority tasks. However, this is not enough for atomicity. The adaptive scheduler of CARTOS (Sec. III-C) waits until enough energy is charged to complete the execution of a peripheral task for this period before scheduling that task. This guarantees that the system does not turn off at least until the completion of the peripheral task, thereby achieving correctness. On the other hand, computational tasks are scheduled preemptively, just like other RTOS, and CARTOS manages their forward progress by JIT checkpointing. To summarize, peripheral and computational tasks in CARTOS have the following characteristics:

### v3

@@ I Introduction
BATTERYLESS systems, often called intermittently-powered devices (IPDs), have the potential to revolutionize a variety of applications in Internet-of-Things (IoT), such as smart healthcare, agriculture, and building monitoring. These systems harvest energy from ambient sources such as sunlight, heat, and radio signals, charge a small energy buffer, and execute intermittently whenever energy is available. This approach addresses critical sustainability challenges in IoT deployments by eliminating battery waste and enabling maintenance-free operation for decades, which has driven the growing industry interest in IPD-based solutions. Especially, IPDs with supercapacitors as their energy buffers offer significant advantages: much longer lifespans than traditional battery-powered systems, deployability in harsh environments (e.g., extreme temperatures) where batteries may fail, smaller form factors, and reduced manufacturing costs [denby2020orbital, choi2022compiler, hester2017future].
The aforementioned benefits, however, come with several technical challenges. First, given the variability in energy availability from ambient sources, IPDs must ensure the forward progress of program execution across power failures. One of the most effective strategies is checkpointing [Chinchilla2018, Idetic2013, Mementos2011, Mayfly2018], a method that involves inserting additional code snippets to store the state of a running program in non-volatile memory (NVM). When power is restored, the system can recover from the latest stored state, ensuring continuous progress. However, such static checkpointing introduces a high overhead and can significantly impact the performance of intermittent systems compared to conventional battery-powered embedded systems. Just-In-Time (JIT) checkpointing [QuickRecall2014, CatNap, Samoyed2019] mitigates this issue by making checkpoints dynamically only when the power goes low. However, it is difficult to be used for tasks with peripheral operations, e.g., sensor access and flash writing, because they cannot be paused and resumed at an arbitrary point, i.e., execution needs to be done atomically.
[…]
@@
CARTOS is implemented by introducing new APIs and components to FreeRTOS, which is one of the most widely used embedded RTOS in real-world applications such as Amazon AWS IoT. However, as the design of CARTOS is not specific to FreeRTOS, it is portable to other RTOS as well. From our experiments on real hardware, CARTOS demonstrated several advantages over the state-of-the-art methods.
Section II provides an overview of previous work and establishes essential background information. In Section II-D, we introduce the proposed framework. Section IV presents the system model, the proposed mixed-preemption scheduling approach, and offers a detailed scheduling analysis. Section V presents the results of our evaluation. Lastly, Section VI summarizes and concludes this paper.
@@ II-A IPD Hardware and Operation Phases
A typical hardware design of IPDs is illustrated in Fig. 1. The harvesting unit collects energy from the surrounding environment. It could be a solar cell, RFID receiver, or piezoelectric transducer. The power management unit (PMU) is responsible for converting and regulating the energy harvested by the energy harvester into a stable voltage that can be used to power the device’s components. It also manages the device’s power consumption to ensure that it does not exceed the available energy. Energy storage is usually a supercapacitor and is responsible for storing the collected electrical energy so that it can be used later when it is needed. The micro-controller unit (MCU) executes the device’s software instructions and controls the device’s various components. Memory stores the program code and data that the MCU uses to operate the device. Most IPDs have at least two types of memory: (i) volatile main memory equipped as part of MCU, e.g., SRAM, and (ii) nonvolatile memory to store execution progress during power loss, e.g., FRAM and MRAM. IPDs also often include sensors and peripherals, such as temperature, humidity, light sensors, as well as wireless communication modules such as Bluetooth LE and LoRa, to monitor the device’s environment, perform certain actions, and report any findings to the user.
@@
    Operation Phase: Once the capacitor attains sufficient voltage, the PMU provides a consistent supply to the device. This enables the MCU to execute tasks. The device continues in this phase until it reaches the predefined power-off threshold (returning to the power-off phase) or low-voltage threshold (moving to the standby phase discussed next) depending on the design. Many prior IPDs [colin2016chain, alpaca2017, Chinchilla2018] use only the power-off threshold for simplicity and operate as in Fig. 2. The MCU starts task execution when the capacitor voltage reaches a specific power-on threshold and continues until the voltage falls below the power-off threshold. Charging can still occur while the system runs tasks in this phase; however, as the discharging rate typically exceeds the charging rate, the capacitor voltage decreases over time.
@@
    Standby Phase: In this phase, energy-intensive components like the MCU, sensors, and radio units are either powered down or transitioned into sleep mode. Only the timekeeping ability, e.g., MCU’s clock unit or external programmable RTC, remains active to maintain the precise time behavior of tasks [MohsenIOTJ2021, Celebi2020, Zygarde, karimi2020energy]. The device remains in this state until the capacitor is sufficiently charged to return to the operation phase.
@@ II-B Forward Progress Mechanisms
Various checkpointing mechanisms have been studied to ensure that the progress of a program on IPDs moves forward with data consistency across power failures. Static checkpointing [Chinchilla2018, Idetic2013, Mementos2011, kortbeek2020time, bhatti2017harvos, kim2023liveness] inserts checkpointing code at the fixed locations, which are determined by either the compiler support or manual programmer efforts. When a checkpoint is reached, the system saves the program’s state and relevant data to non-volatile storage, allowing the program to restart from that point in case of a power failure. Furthermore, new programming models have been studied to create checkpoints as part of the program’s execution flow [Mementos2011, Ink2018, Capybara2018, Colin2018, Mayfly2018, colin2016chain, alpaca2017]. [^1] In the context of IPDs, these approaches are often referred to as task-based intermittent systems. However, it is important to note that “tasks” in these systems differ from our work: they represent short non-preemptible atomic blocks, not-preemptible threads as in conventional RTOS and our work. In such models, the programmer needs to decompose a program logic into a collection of idempotent, atomic blocks with explicit input and output data. Then, the input and output data of each atomic block are recorded in NVM, which serves as checkpoints, e.g., data channels between blocks [colin2016chain, Mayfly2018]. However, this adds a significant burden to programmers to restructure their programs, which is particularly difficult for long-running computations, e.g., algorithm logic. In addition, programmers have to make sure that each atomic block is small enough to run within one charging cycle.
Just-In-Time (JIT) checkpointing [Samoyed2019, QuickRecall2014, Hibernus2015, CatNap], on the other hand, makes checkpoints at runtime by capturing the current state of the system, including MCU register values and volatile data. To do so, the system needs to monitor the remaining energy and initiate JIT checkpointing when the low-voltage threshold is reached. [^2] JIT checkpointing in IPDs [Samoyed2019, QuickRecall2014, Hibernus2015, CatNap], including our work, maintains only the most recent checkpoint, not multiple checkpoint versions, due to inherent storage limitations. This can bring performance benefits over static checkpointing by reducing the frequency of checkpointing and simplifying the programming of long-running computations [Hibernus2015, QuickRecall2014]. Also, the overhead of JIT checkpointing is manageably small due to the use of byte-addressable NVM, e.g., FRAM or MRAM; this means the system does not need to checkpoint data stored in NVM. However, since peripheral operations cannot resume after reboot by recovering just the MCU’s state, JIT checkpointing cannot be used for them [Samoyed2019].
@@
Mementos[Mementos2011]
Chinchilla[Chinchilla2018]
Capybara[Capybara2018]
Quickrecall[QuickRecall2014]
Samoyed[Samoyed2019]
InK[Ink2018]
CatNap[CatNap]
Celebi[Celebi2020]
Rtag[MohsenIOTJ2021]
Karimi et al.[MohenRTCSA2022]
@@ II-C IPD Task Scheduling
Task scheduling for batteryless devices can be categorized into two main types: real-time and reactive scheduling. Real-time scheduling focuses on executing periodic tasks with predictable charging behavior. These approaches aim to meet the deadlines of a specific set of tasks by assuming a certain amount of energy supply[MohsenIOTJ2021, MohenRTCSA2022, Celebi2020]. By doing so, they can guarantee tasks to be scheduled in a predictable manner as long as the assumed energy model holds. On the other hand, reactive scheduling focuses on minimizing the overall response times of event-driven tasks [CatNap, Ink2018]. It starts task execution when the device’s energy reaches the predefined power-on threshold and stops when it drops below the power-off threshold. These approaches are usually more flexible and may perform better with changing environmental conditions than the real-time approaches, but cannot ensure predictability and consistent performance due to their best-effort nature.
@@ II-D Related Work
Among static checkpointing approaches, Mementos [Mementos2011] proposes a method to automatically insert checkpointing code to the program at compile time. To reduce the cost of checkpointing, it also proposes to estimate the remaining energy and perform checkpointing only when the energy goes below a certain threshold. Chinchilla [Chinchilla2018] also takes a similar approach. Capybara [Capybara2018] relies on the programming model support for forward progress, and introduces a hardware-software co-design solution that includes a set of capacitors to handle atomic blocks with different energy requirements. To address the limitations of programming model-based approaches that require the user to divide program logic into a series of atomic blocks, Immortal Threads [yildiz2022immortal] takes a compiler-assisted approach that transforms conventional threaded programs by inserting micro checkpoints. This allows the user to write programs in a typical multithreaded manner and enjoy preemption at the boundaries of checkpoints.
@@
Quickrecall [QuickRecall2014] is a JIT checkpointing system. It detects low-voltage threshold and checkpoints preemptible tasks so that they can be resumed in next power cycle. However, it does not support peripheral tasks and restarts them in next power cycle if the power goes down during their execution. Samoyed [Samoyed2019] addresses this issue by introducing the undo-logging mechanism, which basically adds static checkpoints for peripheral operations. However, this approach is applicable only to the peripherals that “do not have internal non-volatile state and are arbitrarily restartable” [Samoyed2019] because of the idempotency requirement.
@@
Several approaches have been proposed to support peripheral devices in IPDs. Sytare [Sytare] allows users to write additional code to explicitly store peripheral states in NVM. RESTOP [RESTOP] retains peripheral states by recording data transmission between MCU and peripherals, which incurs high overhead. EaseIO [EaseIO] takes a programming model-based approach that allows users to annotate I/O operations and transforms the program code based on this information. These methods offer various trade-offs between programmer effort, system overhead, and generality of peripheral support.
@@
For real-time scheduling on IPDs, Celebi [Celebi2020] proposes a scheduler that optimizes task scheduling based on known charging patterns. However, this method is primarily designed for independent computational tasks and may not handle peripheral tasks correctly because the scheduler does not ensure the atomicity and idempotency of task execution, e.g., the scheduler can stop a task at an arbitrary point. Karimi at al. [MohsenIOTJ2021] introduce a real-time scheduler designed to atomically execute independent periodic tasks with various energy requirements and provide a formal analysis of task schedulability for a fixed charging rate of the system. The same authors [MohenRTCSA2022] present an energy-adaptive scheduler to minimize the age of information. While these approaches contribute to enabling real-time operations on IPDs, none of them are sufficient to handle real-world applications that consist of a chain of peripheral access and long-running computations.
@@
There are also other previous studies that focus on intermittent neural networks [BatterylessNN, Zygarde, MII], multiple energy buffers [Coulombs], energy-aware memory mapping [HybridMem], communication [wardega2022opportunistic], design space exploration [kim2024rapid], and zero-power timekeeping [TimeKeeping]. They contribute to expanding the application scope of IPDs and our work does not conflict with them.
Figure 3: System Framework
@@ III-A Hardware Setup
CARTOS targets an IPD that is equipped with an MCU capable of running a conventional RTOS such as FreeRTOS. The IPD is expected to have both volatile memory (SRAM) and non-volatile memory (NVM). Although we used Ambiq Apollo4 Evaluation board [Apollo4_board] which is equipped with an ARM Cortex-M4 MCU in our implementation, there is no other restriction imposed on MCUs by CARTOS.
@@ III-B Task Chains and Mixed Preemption Model
CARTOS supports periodic applications that consist of a processing chain of tasks, which is a common design pattern found in sensing and intelligent applications [hyunjong_chain, sobhani2023timing, colin2016chain, kim2010decentralized, delgado2021optimal]. Fig 4 illustrates two processing chains, one comprising four tasks and the other with three tasks. Each task $\tau_{i}^{j}$ on a chain $i$ can be either a peripheral or computation task. [^3] Detailed chain and task parameters needed for analysis will be provided in Sec. IV. The chain can start execution when its first task is released (e.g., the periodic sensing task $\tau_{1}^{1}$ on chain 1) and the subsequent tasks are eligible to run only when their preceding task finishes (e.g., the action task $\tau_{1}^{4}$ becomes “ready” after the completion of the processing task $\tau_{1}^{3}$). We assume that the first task of a chain is a periodic task, triggered by either a timer or an external event. Hence, all other tasks of the same chain follow the same periodic pattern with precedence constraints among them. Chains have deadlines to meet. We consider constrained deadlines (deadline $\leq$ period) due to their wide acceptance. A formal model of tasks and chains will follow in Sec. IV.
@@ III-B1 Mixed-Preemption Model
Recall that processing chains can consist of a mixture of peripheral and computation tasks. Peripheral tasks need to be executed atomically, especially when the corresponding hardware maintains internal state [Samoyed2019]. To ensure the atomic execution of peripheral tasks, CARTOS schedules peripheral tasks non-preemptively, i.e., once selected for scheduling, it will continue with no preemption from other higher-priority tasks. However, this is not enough for atomicity. The adaptive scheduler of CARTOS (Sec. III-C) waits until enough energy is charged to complete the execution of a peripheral task for this period before scheduling that task. This guarantees that the system does not turn off at least until the completion of the peripheral task, thereby achieving correctness. On the other hand, computational tasks are scheduled preemptively, just like other RTOS, and CARTOS manages their forward progress by JIT checkpointing. To summarize, peripheral and computational tasks in CARTOS have the following characteristics:
@@
III-B2 Memory Model
Peripheral (atomic) tasks use SRAM for data and stack segments since their memory states need to be initialized upon reboot. In contrast, computational (non-atomic) tasks can utilize NVM for data segments to preserve data across power failures and reduce JIT checkpointing overhead. Stack areas of non-atomic tasks could also be allocated in NVM, but our implementation allocates them in SRAM for simplicity and because their memory footprint is relatively small. OS-level data structures are maintained in SRAM, with those requiring persistency (e.g., timer events, contexts of non-atomic tasks, and profiling data) being selectively checkpointed. We assume static memory allocation only, as dynamic memory allocation is typically prohibited in real-time systems due to timing unpredictability, and is not considered in our framework.